OpenStudy (anonymous):

How do I "check" this by substituting each value x=2 & x=4 into log5(x+3)+log5(x-1)...

OpenStudy (campbell_st):

\[\log _{5} (2 + 3) = 1\] and \[\log _{5} (2 -1) = 0\] so when x = 2 the answer would be 1 for a base 5 log system

OpenStudy (campbell_st):

same process for x = 4

OpenStudy (anonymous):

so when u say same process for x=4 do i put 4 into the equation as well

OpenStudy (campbell_st):

yes... repeat the process for x = 4.... its log(4+3) + log (4-1)

OpenStudy (anonymous):

so i should come up with 1 and 0 again in order to come out right

OpenStudy (campbell_st):

no you have log 7 and log 3 The reason the 1st part was so easy in base 5 logs, 5 = 5^1 so then answer is 1 and 5^0 = 1 so the answer is 0...

OpenStudy (anonymous):

o ok so there is no substituting you just finish it out

OpenStudy (anonymous):

so when u have log7(4+3)=1 and log3(4-1)=1

OpenStudy (campbell_st):

well the only other suggestion is log(x+3) + log(x-1) is the same as log{(x+3)(x-1)} using log laws.... so it could be log (7 x 3) = log 21

OpenStudy (anonymous):

ok???

OpenStudy (campbell_st):

lol... sorry... just they are funny numbers for base 5 logs.... change of base is an option if you need a number answer

OpenStudy (anonymous):

ok so earlier when i put log7(4+3)=1 & log3(4-1)=1 would that be right